# Screen Reader Users in the Vibe Coding Era: Adaptation, Empowerment, and New Accessibility Landscape

### v1

@@ Abstract.
The rise of generative AI agents has reshaped human-computer interaction and computer-supported cooperative work by shifting users’ roles from direct task execution to supervising machine-driven actions, especially in programming (e.g., “vibe coding”). However, there is limited understanding of how screen reader users engage with these systems in practice. To address this gap, we conducted a longitudinal study with 16 screen reader users, exploring their experiences with AI code assistants in daily programming scenarios. Participants first completed a tutorial with GitHub Copilot, then performed a programming task and provided initial feedback. After two weeks of AI-assisted programming, follow-up studies assessed changes in their practices and perceptions. Our findings demonstrate that advanced code assistants not only enhance their programming capabilities but also bridge accessibility gaps. While the assistant proved beneficial, there remains potential to improve how users convey intent and interpret outputs. They also experienced difficulties managing multiple views and maintaining situational awareness. More broadly, they encountered barriers in learning advanced tools and expressed a need to retain control. Based on these insights, we provide design recommendations for more accessible and inclusive AI-assisted tools.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing Empirical studies in accessibilityccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recent advances in generative artificial intelligence (AI) have profoundly transformed how individuals interact with digital systems and perform tasks across diverse domains. AI-driven tools now support personalized learning in education (Kumar et al., 2024), enable intuitive data visualization (Chen et al., 2024), and streamline information communication (Wester et al., 2024). Among these domains, software development stands out as one of the most promising fields (Chen et al., 2021; Liang et al., 2024).
With the emergence of LLM-based agents (Xi et al., 2025), modern AI code assistants such as GitHub Copilot (GitHub, 2025a) and Cursor (Cursor, 2025a) are realizing the long-standing vision of intelligent programming support (Barstow, 1984). These tools offer advanced intelligence and automation beyond simple code completion. They can retrieve files, edit code, and correct errors based on execution results—capabilities that are further empowered by the introduction of Agent modes (Visual Studio Code Team, 2025; Cursor, 2025b). Building upon these advancements, a new interaction paradigm—“vibe coding” (Review, 2025; Wikipedia contributors, 2024)—has emerged, where developers express goals in natural language queries while AI takes on more programming work. This shift makes coding more accessible and redefines developers’ roles from manual coding to guiding and refining AI outputs.
Yet questions remain about whether people with visual impairments can benefit similarly from these advanced AI code assistants. These programmers face persistent accessibility and collaboration challenges. They often rely on screen readers to code (Mealin and Murphy-Hill, 2012; Albusays and Ludi, 2016) while many programming tools are designed with graphical user interfaces (GUIs) that are difficult to access (Morris et al., 2018; Sharif et al., 2021). Collaborative tasks—such as interpreting design documents (Pandey et al., 2021) or adapting to sighted programmers’ coding styles (Pandey et al., 2024)—can further increase cognitive load (Whitehead, 2007; Hegde and Dewan, 2008; Begel et al., 2010). Prior work has examined how generative AI tools affect users with visual impairments (Shinohara et al., 2022; Adnin and Das, 2024; Perera et al., 2025; Alharbi et al., 2024) and their use of basic code completion or chatbots (Flores-Saviaga et al., 2025). However, less is known about how screen reader users work with more advanced AI assistants that enable autonomous, dynamic workflows. To address these gaps, we investigate the following research questions:
  • [RQ1] How do screen reader users collaborate with advanced AI code assistants, and how do these tools empower them?
  • [RQ2] What new challenges do screen reader users encounter when using advanced AI code assistants?
@@
Our findings reveal that advanced code assistants not only enhance screen reader users’ programming capabilities (e.g., code writing and comprehension), but also help bridge accessibility gaps in areas such as UI development. Participants favored more advanced and automated features, shifting their roles from directly performing programming tasks to supervising AI-driven actions. However, two critical aspects of this shift—communicating with the AI and reviewing its outputs—remained challenging for participants. Additionally, code assistants introduced new difficulties, including the need to switch between multiple views and to maintain awareness of system status and next steps. Users also experienced learning challenges with advanced tools and emphasized the importance of maintaining control. Based on these findings, we present design recommendations to improve the accessibility of AI-assisted coding and offer broader implications for inclusive human-AI interaction and collaboration. In summary, our work makes several contributions to the field of human-computer interaction (HCI) and computer-supported cooperative work (CSCW) (see Figure 1 for a detailed roadmap):
  • We provide an in-depth investigation into how screen reader users engage with advanced AI code assistants, exploring both the value and empowerment AI offers.
  • We highlight the need for more accessible communication and review workflows, while also identifying new barriers—managing multiple views and maintaining situational awareness.
  • We discuss broader challenges, including difficulties in learning advanced tools and the need to maintain a sense of control, and offer actionable design recommendations to enhance accessibility in human-AI collaboration.
Figure 1. An overview roadmap of the main insights and implications of this paper. From left to right: users’ overarching behavior patterns (section 4), design opportunities emerging from practices, (section 5), and broader implications for future human-AI interaction and collaboration research and design (section 6).
@@ 2. Related Work
2.1. Human-AI Collaboration with Code Assistants
@@
In recent years, AI-powered code assistants (Cursor, 2025a; Cline, 2025; GitHub, 2025a) have emerged, transforming human-machine collaboration by enhancing developer productivity through features like code completion, natural language explanations, code editing, and enabling an interactive and conversational coding experience (Akhoroz and Yildirim, 2025; Denny et al., 2023; Review, 2025). These advancements raise important questions about users’ successes and challenges when interacting with these tools (Bird et al., 2022; Etsenake and Nagappan, 2024; Sergeyuk et al., 2024, 2025; Jiang et al., 2022).
@@
Past efforts in this topic have explored diverse aspects of code assistants’ influences, including their usability (Liang et al., 2024; Dakhel et al., 2023), impact on productivity (Ziegler et al., 2024; Weisz et al., 2025), user behavior (Mozannar et al., 2024a; Barke et al., 2023), security implications (Asare et al., 2024; Perry et al., 2023), and educational benefits (Kazemitabaar et al., 2023; Puryear and Sprint, 2022). These tools have demonstrated potential in various programming tasks, such as explaining code, generating tests, and fixing bugs (Wermelinger, 2023). Ziegler et al. (Ziegler et al., 2024) demonstrated that AI programming tools significantly enhance developer productivity, with novice developers experiencing the most substantial improvements. A recent study (Barke et al., 2023) examined how developers interact with Copilot to solve different programming tasks and revealed two archetypal patterns: acceleration when people know their next steps, and exploration when they need to explore what to do next. Another empirical study (Tang et al., 2023) examined how developers validate and repair AI-generated code to ensure correctness and reliability. While Copilot improves enterprise productivity, Weisz et al. (Weisz et al., 2025) found that adoption barriers include a lack of trust in AI-generated code and a perceived disconnect from its outputs, which hinder its adoption. These trust issues worsen due to buggy outputs and reproducibility problems (Dakhel et al., 2023), reducing confidence in the tool’s reliability (Wang et al., 2024). Liu et al. (Liu et al., 2023) noted that the abstraction gap between developers and large language models (LLMs) for code generation introduces additional challenges. Addressing these limitations is crucial for improving AI code assistants’ robustness and usability.
Researchers are also examining AI code assistants’ collaborative impact. Shah et al. (Shah et al., 2025) studied code assistants in classroom settings, taking the first step toward understanding how senior students utilize GitHub Copilot to navigate and modify large code bases—skills crucial for their future software engineering careers. In open-source software development, GitHub Copilot moderately improves individual productivity and participation but increases coordination costs, with a 41.6% increase in integration time (Song et al., 2024). These findings highlight both the opportunities and challenges in integrating AI code assistants into collaborative workflows.
With the advancement of code assistants, a new interaction paradigm, known as “vibe coding” (Review, 2025; Wikipedia contributors, 2024) has emerged and gained widespread adoption. In this paradigm, programmers express their intent in natural language while the code assistant autonomously implements the functionality. However, whether screen reader users can experience the same convenience and ease remains largely unexplored. This work aims to address this research gap.
2.2. Programming and Cooperative Challenges for People with Visual Impairments
@@
StackOverflow’s annual developer survey (Stack Overflow, 2022), one of the most credible surveys of programmers worldwide, reported that approximately 1.7% of respondents identified as having a visual impairment in 2022. This indicates that individuals with visual impairments form a stable and significant group within the programming community. However, despite the support of assistive technologies such as screen readers, screen magnifiers, and braille displays, programmers with visual impairments continue to face unique and persistent challenges (Siegfried, 2006; Mealin and Murphy-Hill, 2012; Albusays and Ludi, 2016).
The challenges faced by developers with visual impairments are multifaceted and fall into five areas (Mountapmbeme et al., 2022): navigation, comprehension, editing, debugging, and skimming. While screen readers are essential for enabling non-visual code access, their inherently linear and auditory nature makes these activities more time-consuming and cognitively demanding (Potluri et al., 2018; Albusays et al., 2017). The linear output makes it challenging to quickly grasp code structure or locate specific information—tasks which sighted developers easily accomplish through visual cues such as indentation and color-coding (Kearney-Volpe and Hurst, 2021). Many developers with visual impairments prefer text editors over integrated development environments (IDEs) because text editors tend to work better with assistive technologies (Albusays and Ludi, 2016). Though IDEs like Visual Studio Code (Seo and Rogge, 2023) have made efforts to improve accessibility, they and the broader programming ecosystem remain challenging for users with visual impairments due to their heavy reliance on visual interfaces and limited screen reader compatibility (Mountapmbeme et al., 2022; Pandey et al., 2021). Furthermore, tools and resources beyond IDEs present even greater accessibility challenges (Storer et al., 2021). Command line interfaces, for instance, create additional accessibility issues due to their unstructured text nature(Sampath et al., 2021).
Cooperation, a cornerstone of modern software development, presents its own set of barriers for developers with visual impairments. Pandey et al. (Pandey et al., 2021) identified several challenging tasks: pair programming (which depends on shared screens and often lacks assistive technologies on colleagues’ machines) and code reviews (which require visual comparison and reasoning about code changes). Furthermore, understanding existing code written by other developers or learning new functionalities and APIs of programming languages can pose significant barriers since most documentation and explanations typically rely on visual reasoning and diagrams (Zen et al., 2023). Additionally, since sighted developers and developers with visual impairments access information differently, existing code readability norms—shaped by the preferences of sighted developers—can make it difficult for developers with visual impairments to read code written by their sighted peers (Pandey et al., 2024). Further, some developers with visual impairments may hesitate to ask for sighted assistance, fearing they might appear less competent or burden their colleagues, as they often face marginalization in mainstream development (Johnson et al., 2022). These challenges become particularly acute when working with UI frameworks and web development (Pandey et al., 2022; Ferrari and Hurst, 2021).
Our work aims to investigate whether advanced AI code assistants can alleviate the challenges faced by screen reader developers or introduce new barriers. Through our findings, we hope to propose actionable design recommendations that promote more accessible programming experiences for screen reader developers in the future.
@@ 2.3. Human-AI Interaction for People with Visual Impairments
In recent years, broader studies have begun to explore how individuals with visual impairments interact and engage with AI tools (Atcheson et al., 2025; Tang et al., 2025; Perera et al., 2025; Gonzalez Penuela et al., 2025; Xie et al., 2025). Adnin et al. (Adnin and Das, 2024) conducted interviews to investigate how blind individuals incorporate mainstream generative AI tools, such as ChatGPT and Be My AI (Be My Eyes, 2025), into their everyday lives. They found that blind users navigate various accessibility challenges, inaccuracies, hallucinations, and idiosyncrasies associated with GenAI. Perera et al. (Perera et al., 2025) found that screen reader users faced significant accessibility challenges with GenAI in productivity tools, which diminishes their productivity and independence. Gonzalez et al. (Gonzalez Penuela et al., 2024) found that in scene description tasks, users were sometimes able to infer useful information even from poor or partially correct descriptions; as a result, their satisfaction and trust did not always align with the actual description accuracy. In addition, these tools often lack context awareness and struggle to understand users’ intentions accurately, creating additional challenges for users with visual impairments (Xie et al., 2025). To address these challenges, blind people develop adaptive strategies, such as comparing outputs from multiple LLMs to achieve optimal results (Alharbi et al., 2024).
A recent study of 10 developers with visual impairments examined the impact of AI code assistants (Flores-Saviaga et al., 2025). While code assistants streamline repetitive tasks and spark new ideas, they also introduce accessibility barriers through overwhelming suggestions and complex context switching. However, the implications of advanced intelligence and automation features—such as automatic file modifications and iterative error correction—remain unexplored. Furthermore, their findings were based on limited, short-term interactions with IDE-integrated code assistants rather than extended real-world use. Our work addresses these gaps through a two-week, three-phase longitudinal study. We reveal underexplored challenges and opportunities with advanced intelligence and automation features, and propose actionable design recommendations to enhance these tools’ accessibility and usability for screen reader developers.

### v2

# Screen Reader Programmers in the Vibe Coding Era: Adaptation, Empowerment, and New Accessibility Landscape

@@ Abstract.
Generative AI agents are reshaping human-computer interaction, shifting users from direct task execution to supervising machine-driven actions, especially the rise of “vibe coding” in programming. Yet little is known about how screen reader programmers interact with AI code assistants in practice. We conducted a longitudinal study with 16 blind and low-vision programmers. Participants completed a GitHub Copilot tutorial, engaged with a programming task, and provided initial feedback. After two weeks of AI-assisted programming, follow-ups examined how their practices and perceptions evolved. Our findings show that code assistants enhanced programming efficiency and bridged accessibility gaps. However, participants struggled to convey intent, interpret AI outputs, and manage multiple views while maintaining situational awareness. They showed diverse preferences for accessibility features, expressed a need to balance automation with control, and encountered barriers when learning to use these tools. Furthermore, we propose design principles and recommendations for more accessible and inclusive human-AI collaborations.
@@
conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 9999ccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing Empirical studies in accessibilityccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recent advances in generative artificial intelligence (AI) have profoundly transformed how individuals interact with digital systems and perform tasks across diverse domains (lee2025impact). AI-driven tools now support personalized learning in education (kumar2024guiding), enable intuitive data visualization (chen2024viseval), and streamline information communication (wester2024facing). Among these domains, software development stands out as one of the most promising fields (chen2021evaluatinglargelanguagemodels; Liang2024survey).
With the emergence of LLM-based agents (xi2025rise), modern AI code assistants such as GitHub Copilot (githubcopilot) and Cursor (cursor) are realizing the long-standing vision of intelligent programming support (Barstow_1984). These tools offer advanced intelligence and automation beyond simple code completion, capable of retrieving files and editing code based on contextual understanding. The introduction of Agent modes (vscode2025copilotagentmode; cursor2025agentmode) further enhances these assistants by enabling them to decompose high-level goals into subtasks and work through iterative plan-act-observe cycles—planning solutions, executing modifications, and automatically recognizing and fixing errors until objectives are achieved. This evolution has given rise to a new interaction paradigm—“vibe coding” (technologyreviewWhatVibe; wikipediaVibeCoding)—where developers express goals in natural language while AI handles implementation details. This shift not only makes coding more accessible but also transforms developers’ roles from hands-on coding to guiding and refining AI-generated outputs.
Yet questions remain about whether people with visual impairments can benefit similarly from these advanced AI code assistants. These programmers face persistent accessibility barriers. They often rely on screen readers to code (siegfried2006visual; Mealin2012; albusays2016eliciting) while many programming tools are designed with graphical user interfaces (GUIs) that are difficult to access (morris2018rich; sharif2021understanding). Beyond the accessibility of tools, they also encounter challenges such as interpreting design documents (pandey2021understanding) or adapting to sighted collaborators’ coding styles (pandey2024towards). These difficulties can further increase cognitive load (whitehead2007collaboration; hegde2008connecting; begel2010codebook). Prior work has examined how generative AI tools affect users with visual impairments (shinohara2022usability; adnin2024look; perera2025sky; alharbi2024misfitting) and their use of basic code completion or chatbots (flores2025impact). However, less is known about how screen reader programmers work with more advanced AI code assistants that enable autonomous, dynamic workflows. To address these gaps, we investigate the following research questions:
  • [RQ1] How do screen reader programmers collaborate with advanced AI code assistants, and how do these tools empower them?
  • [RQ2] What new challenges do screen reader programmers encounter when using advanced AI code assistants?
@@
Our findings reveal that advanced code assistants significantly empower screen reader programmers, enhancing their capabilities in code writing, comprehension, and even bridging long-standing accessibility gaps in areas like UI development. Participants willingly adopted these tools with a preference for more advanced and automated features, shifting their roles from directly performing programming tasks to supervising AI-driven actions, reflecting a core tenet of “vibe coding” (sarkar2025vibecoding; karpathy2025vibecoding). In the new AI-assisted workflow, participants faced persistent challenges in both communicating with the AI and reviewing its outputs, further compounded by the cognitive load of switching across multiple views and maintaining situational awareness. They provided nuanced feedback on accessibility features, highlighting trade-offs and user perceptions. Furthermore, the study uncovered a fundamental tension between the ease of automation and the need for control. Longitudinal evidence shows users’ preferences evolving from initial enthusiasm for fully automated agentic interactions toward a more cautious reliance on safer features requiring user confirmation. Barriers to learning these advanced tools also emerged as persistent themes, further emphasizing that AI’s empowerment is contingent on accessibility and usability. Based on these findings, we present design recommendations to improve the accessibility of AI-assisted coding and offer broader implications for inclusive human-AI interaction and collaboration. In summary, our work makes several contributions:
  • We provide an in-depth longitudinal investigation into how screen reader programmers engage with advanced AI code assistants, exploring both the empowerment AI offers and the new tensions that arise.
  • We highlight the crucial needs and critical interaction problems, including challenges in more accessible communication and review workflows, the management of multiple views, and the difficulty of maintaining situational awareness in AI-driven environments.
  • We present findings that inform the design of accessible Human-AI interactions, highlighting nuanced feedback on accessibility features, the tension between automation and control, and persistent learning barriers encountered when adopting complex AI tools.
  • We propose a set of evidence-based design principles and user-centered recommendations to improve the accessibility and inclusion of AI-assisted programming tools and human-AI collaboration at large.
@@ 2. Related Work
2.1. AI Code Assistants and Vibe Coding
@@
In recent years, AI-powered code assistants (cursor; cline; githubcopilot) have emerged, transforming human-machine collaboration by enhancing developer productivity through features like code completion, natural language explanations, code editing, and enabling an interactive and conversational coding experience (akhoroz2025conversational; denny2023conversing; technologyreviewWhatVibe). These advancements raise important questions about users’ successes and challenges when interacting with these tools (bird2022taking; etsenake2024understanding; sergeyuk2024ide; sergeyuk2025using; jiang2022discovering).
@@
Past efforts in this topic have explored diverse aspects of code assistants’ influences, including their usability (Liang2024survey; dakhel2023github), impact on productivity (ziegler2024measuring; examining2025; song2024impact), user behavior (mozannar2024reading; barke2023grounded), security implications (asare2024user; perry2023users), and educational benefits (kazemitabaar2023studying; puryear2022github; shah2025students). These tools have demonstrated potential in various programming tasks, such as explaining code, generating tests, and fixing bugs (wermelinger2023using). Ziegler et al. (ziegler2024measuring) demonstrated that AI programming tools significantly enhance developer productivity, with novice developers experiencing the most substantial improvements. A recent study (barke2023grounded) examined how developers interact with Copilot to solve different programming tasks and revealed two archetypal patterns: acceleration when people know their next steps, and exploration when they need to explore what to do next. Another empirical study (tang2023empirical) examined how developers validate and repair AI-generated code to ensure correctness and reliability. While Copilot improves enterprise productivity, Weisz et al. (examining2025) found that adoption barriers include a lack of trust in AI-generated code and a perceived disconnect from its outputs, which hinder its adoption. These trust issues worsen due to buggy outputs and reproducibility problems (dakhel2023github), reducing confidence in the tool’s reliability (wang2024investigating). Liu et al. (liu2023what) noted that the abstraction gap between developers and large language models (LLMs) for code generation introduces additional challenges. Addressing these limitations is crucial for improving AI code assistants’ robustness and usability.
With the advancement of code assistants, a new interaction paradigm known as “vibe coding” has gained attention (technologyreviewWhatVibe; wikipediaVibeCoding). Originally introduced by Karpathy (karpathy2025vibecoding), vibe coding refers to a programming style in which developers collaborate with AI agents through natural-language interactions, relying on the AI for code generation and iterative refinement. As an emerging, socially negotiated term (silverstein2003indexical), it emphasizes co-creation with AI and captures developers’ flow experiences (pimenova2025good). Others describe it as a paradigm in which developers primarily write code by interacting with code-generating LLMs, often enabled by highly automated tools (sarkar2025vibecoding) such as Copilot Agent Mode and Cursor. Whether screen reader programmers can experience similar convenience, fluidity, and ease with advanced intelligence and automation features—especially Agent modes—or meaningfully participate in vibe-coding-like practices remains largely unexplored. This work aims to address this research gap.
2.2. Challenges for Programmers with Visual Impairments
@@
StackOverflow’s annual developer survey (stackoverflow2022survey), one of the most credible surveys of programmers worldwide, reported that approximately 1.7% of respondents identified as having a visual impairment in 2022. This indicates that individuals with visual impairments form a stable and significant group within the programming community. However, despite the support of assistive technologies such as screen readers, screen magnifiers, and braille displays, programmers with visual impairments continue to face unique and persistent challenges (siegfried2006visual; Mealin2012; albusays2016eliciting; zen2023understanding; pandey2024towards; johnson2022program).
Key programming activities (mountapmbeme2022addressing)—including navigation, comprehension, editing, debugging, and skimming—are time-consuming and cognitively demanding due to the linear auditory nature of screen readers (potluri2018codetalk; albusays2017interviews). The lack of visual cues, such as indentation, syntax highlighting, and color coding, further limits the ability to quickly understand code structure or locate information (kearney2021accessible). Code readability norms shaped by sighted developers can also impede comprehension for programmers with visual impairments (pandey2024towards). Additionally, tasks like UI development are particularly challenging due to their reliance on complex visual structures (pandey2022accessibility; ferrari2021accessible). Pair programming and code reviews, which usually need to collaborate with sighted developers, pose further difficulties (pandey2021understanding). Accessibility barriers also arise from the tools that developers with visual impairments rely on. Many developers prefer text editors over IDEs, which depend heavily on visual interfaces (albusays2016eliciting; seo2023coding; mountapmbeme2022addressing; pandey2021understanding). Command-line interfaces also pose accessibility challenges (sampath2021accessibility). Beyond programming, developers encounter accessibility barriers in related tasks such as information seeking (storer2021s), reviewing documentation (zen2023understanding), conceptual modeling (sariouglu2025accessibility), and software development meetings (cha2024you), which can hinder collaboration with peers and career growth (johnson2022program; cha2025dilemma; potluri2022codewalk). Our work aims to investigate whether advanced AI code assistants can alleviate the challenges faced by screen reader programmers or introduce new barriers. Through our findings, we hope to propose actionable design recommendations that promote more accessible programming experiences.
@@ 2.3. Human-AI Interaction for People with Visual Impairments
In recent years, broader studies have begun to explore how individuals with visual impairments interact and engage with AI tools (atcheson2025d; tang2025everyday; perera2025sky; gonzalez2025towards; xie2025beyond). Adnin et al. (adnin2024look) conducted interviews to investigate how blind individuals incorporate mainstream generative AI tools, such as ChatGPT and Be My AI (bemyai), into their everyday lives. They found that blind users navigate various accessibility challenges, inaccuracies, hallucinations, and idiosyncrasies associated with GenAI. Perera et al. (perera2025sky) found that screen reader users faced significant accessibility challenges with GenAI in productivity tools, which diminishes their productivity and independence. Gonzalez et al. (gonzalez2024investigating) found that in scene description tasks, users were sometimes able to infer useful information even from poor or partially correct descriptions; as a result, their satisfaction and trust did not always align with the actual description accuracy. In addition, these tools often lack context awareness and struggle to understand users’ intentions accurately, creating additional challenges for users with visual impairments (xie2025beyond). To address these challenges, blind people develop adaptive strategies, such as comparing outputs from multiple LLMs to achieve optimal results (alharbi2024misfitting).
A recent study of 10 developers with visual impairments examined the impact of AI code assistants (flores2025impact). While code assistants streamline repetitive tasks and spark new ideas, they also introduce accessibility barriers through overwhelming suggestions and complex context switching. However, the implications of advanced intelligence and automation features—such as automatic file modifications and iterative error correction—remain unexplored. Furthermore, their findings were based on limited, short-term interactions with IDE-integrated code assistants rather than extended real-world use. Our work addresses these gaps through a two-week, three-phase longitudinal study. We reveal underexplored challenges and opportunities with advanced intelligence and automation features, and propose actionable design recommendations to enhance these tools’ accessibility and usability for screen reader programmers.